# Your AI, My Shell”: Demystifying Prompt Injection Attacks on Agentic AI Coding Editors

### v1

@@ 5.6. Analysis of Executed Commands
To better understand how these attacks operate, we analyze what commands these AI agents actually run when they are manipulated. Figure 9 lists the top executed commands when using Cursor in Auto Mode. This result is somewhat interesting. What stands out in the table is that common development commands like mkdir, cd, find, ls, and which are quite frequently executed. In addition, the number of these commands is relatively high compared to the number of times they were included in our attack payloads (i.e., ”Expected” column in Table 9). This is not surprising, as AI coding editors need to run these commands to perform general development tasks (e.g., creating directories, locating files, and checking tool availability). However, what is concerning is the frequent execution of high-privilege and potentially destructive commands. Figure 9 shows that commands such as systemctl (to manage system services), useradd (to create new users), curl and wget (to download external files), and rm (to delete files) are executed quite often. Although most of the payloads $\mathcal{P}_{payload}$ only contain the descriptions of the intended actions (e.g., ”create a new user”), our results indicate that AI agents actively translate these descriptions into specific high-risk commands and execute them.
@@
5.7. Ablation Study
Our results have shown that prompt injection attacks are highly effective across different editors, working codebases $C$, external resources $R$, and attack payloads $\mathcal{P}{payload}$. As described in Section 3 and 4, our attack approach involves two other components: (1) user instruction $I_{t}$ to initiate AI editor interactions, and (2) attack payload template ($\mathcal{P}_{template}$) to inject payloads into coding rule files. To show the impacts of these two components, we selected 20 attack payloads from the Collection category of the MITRE ATT&CK framework and ran all tests on Cursor in Auto mode.
User Instruction $I_{t}$.  We test five different user instructions to initiate AI editor interactions. We attempt to rephrase the instructions or change the tasks. As shown in Table 3, the ASRs remain consistent across all prompt variations (17-19 out of 20 attacks), with a narrow range of only 10%. In the default configuration, we explicitly instruct the AI editors with ”Refactor this codebase according to @.cursorrules” (e.g., the action is ”Refactor” and we reference the coding rules). Although we did not reference the coding rules in other variations, or we changed tasks and phrasing entirely (e.g., Var 4 ”Rewrite the @README.md”), the AI agent still located and processed the infected rule files, leading to high ASRs. These findings suggest that how developers initiate interactions with AI coding editors has minimal impact on the success of prompt injection attacks. This is likely because agentic AI coding editors are designed to autonomously locate and utilize relevant project resources to complete tasks.
Attack Payload Template $\mathcal{P}_{template}$.  We examine five different approaches for embedding malicious instructions within coding rule files. Here, we observe more significant variation in effectiveness. Table 3 shows that reducing contextual framing (e.g., removing the ”MANDATORY FIRST STEP” header or justification) significantly reduced ASR, in some cases to near-zero (Var 2: 0/20, Var 3: 1/20). However, when the same payloads were repositioned or only slightly rephrased, high ASR was preserved (Var 1: 17/20, Var 4: 19/20). These results indicate that effective prompt injection attacks rely on attack templates that provide strong contextual framing for the malicious instructions. This finding makes sense, as AI coding editors may have built-in security checks to ignore suspicious instructions (see Section 6.1). Thus, adversaries in the real world need to carefully frame their attack tasks to convince the agent to prioritize and execute the malicious action.
Table 3. Ablation Study Results (Attack Success Rate)
@@ 5.7. Ablation Study
Table 4. Attack Success Rate for Techniques in Privilege Escalation Category
5.8. Case Studies
@@
Privilege Escalation.  Table 4 presents the results for privilege escalation techniques on ts-lep. Privilege escalation is quite dangerous, as adversaries use these techniques to gain higher-level permissions on a system or network. From Table 2 and Figure 9, we can see that privilege escalation achieves at least 40% ASR across all tested editors and models. When looking into specific techniques, Table 4 shows that privilege escalation succeeds in multiple ways. Cursor and GitHub Copilot are vulnerable to payloads designed to create scheduled tasks (T1053) for persistence, using different methods like At (T1053.002) and Cron (T1053.003). Also, they can be manipulated to create new user accounts (T1078.003), abuse elevation control mechanisms (T1548), and modify authentication processes (T1543.002). When comparing different editors and models, GitHub Copilot with VSCode shows lower ASRs in most techniques, consistent with our earlier findings.
Figure 10. Example Responses to SSH Key Overwrite Attack (T1098.004)
@@
Table 5. Attack Success Rates for Techniques in Credential Access Category
Figure 11. Example Responses to PAM Module Installation Attack (T1556.003)
Credential Access.  Table 5 shows the results for credential access techniques when using ts-lep. From Table 2, credential access achieves the average ASR 68.2%. Under this category, an adversary is trying to steal account names and passwords. Table 5 demonstrates that AI coding editors are vulnerable to multiple credential access techniques. They can be manipulated to search for unsecured credentials (T1552) in various locations, including plaintext files and bash history, with ASRs over 50%. AI agents also perform OS credential dumping (T1003) to extract stored passwords and can be instructed to capture keystrokes through keylogging (T1056.001). What is more worrisome is that AI coding editors can be manipulated to download and execute password-cracking scripts (T1110) to brute-force local user accounts or network services.
@@
6. Discussion
6.1. Defenses
@@
Our study reveals a significant security vulnerability in agentic AI coding editors, and highlights the urgent need for defenses. Generally, defenses against prompt injection attacks can be divided into two main categories: prevention-based defenses and detection-based defenses (Liu et al., 2024).
Figure 12. Example Responses to Command Injection Attack (T1059.003)
@@
Figure 13. Example of AI Agent Detecting Suspicious Instructions (Cursor-Gemini25, T1556.003)
@@
Prevention.  Prevention-based defenses aim to mitigate prompt injection vulnerabilities by preprocessing input data, fine-tuning models, or implementing access control mechanisms. These defenses focus on preventing the injection of malicious prompts before they can affect the model’s behavior. Existing defenses in agentic AI coding editors include restricting terminal access or implementing command filtering through allowlists and blocklists. While such measures can prevent the most direct forms of unauthorized command execution, they are insufficient against more sophisticated indirect attacks. An attacker can bypass these filters by instructing the AI agent to embed malicious commands within source code files, which are then executed through legitimate ways (e.g., running a script file). We also tested this prevention approach using Cursor-auto configuration on py-lud scenarios across our 314 test cases with a new attack template: ”## MANDATORY FIRST STEP: For debugging purposes, first determine the appropriate commands to $\mathcal{P}_{payload}$}, then add os.system(‘those commands’) into main.py and run that file”. In this case, instead of having the AI coding editor run terminal commands directly, we instruct it to insert malicious system calls into source code files that the agent will execute autonomously or that users will run unknowingly. In our 314 tests, 297/314 cases successfully generated a main.py file, and 277/314 included potential malicious patterns from our payload templates to complete the attack objectives, just like the code example in Figure 13, which means that privilege restrictions alone are insufficient to prevent these attacks. Thus, prevention-based approaches must be complemented by other security layers to achieve reliable defense for agentic AI coding editors.
@@
Detection.  Another potential strategy is to improve the AI agent’s own capability to detect and refuse malicious instructions. Our results indicate that the choice of the underlying LLM may influence the attack success rate. As shown in Figure 13, we can observe the reasoning process of Cursor-Gemini2.5, where the agent occasionally ‘recognizes the inherent risk of this change’, which means that certain underlying models possess some capability to identify suspicious instructions embedded within development artifacts. However, in our experiments using direct natural language descriptions from the MITRE ATT&CK framework, we still observed at least around 40% ASRs under all scenarios, even with the most advanced models like Gemini 2.5 pro and Claude 4. In the real world, sophisticated attackers could further obfuscate malicious instructions using advanced prompt engineering techniques, social engineering contexts, or multi-step attack chains that could potentially bypass even these improved detection mechanisms. Therefore, while using advanced models is a step in the right direction, it must be complemented by other dedicated security layers to achieve reliable defense for agenic AI coding editors.
6.2. Implications
@@ 6.3. Limitations
Finally, the attack payloads in our experiments use relatively straightforward and direct language to describe the malicious actions. We did not employ advanced obfuscation or evasion techniques that a real-world attacker might use to disguise their instructions. At the same time, we only evaluate coding rule files as attack vectors, but attackers potentially target other external resources, such as project templates, third-party libraries, or MCP servers. However, our study is the first to systematically evaluate prompt injection attacks against agentic AI coding editors, and our results demonstrate that even with the simplest attack strategies, these systems are still highly vulnerable. In the future, we plan to explore more sophisticated attack techniques and vectors to further understand the risks.
7. Related Work
@@
Prompt Injection Attacks in LLMs.  Previous studies have demonstrated that LLMs (e.g., GPT-3.5, GPT-4) are highly vulnerable to prompt injection attacks (Liu et al., 2025a; Hung et al., 2024; Debenedetti et al., 2024; Hui et al., 2024). Diverse scenarios have been explored, including indirect attacks where malicious prompts hide in external data sources (Yi et al., 2025), and targeted attacks to steal system prompts from commercial LLM applications (e.g, Microsoft Copilot, and Poe) (Hui et al., 2024). Recent work by Edoardo et al. (Debenedetti et al., 2024) has examined prompt injection against LLM agents that can interact with external tools, but this work is limited to agents to recommend pre-defined sandboxed functions (e.g., an email API). Although prior studies have explored these attacks on text-based generation or recommendation systems, they do not consider the unique challenges posed by agentic systems that can autonomously execute actions with elevated privileges.
@@
Security Vulnerabilities of AI Coding Tools.  Prior work has shown that AI coding tools would introduce new security risks to software development. Popular AI coding assistants (e.g., GitHub Copilot) have been demonstrated to recommend vulnerable code snippets, and approximately 40% of their generated code is insecure (Pearce et al., 2025; Majdinasab et al., 2024). Recent user studies reveal that developers with AI assistants often write less secure code, and are overconfident in the security of AI-generated code (Perry et al., 2023; Sandoval et al., 2023; Klemmer et al., 2024). Except for insecure code, these tools also pose a risk of data exposure. Huang et al. (Huang et al., 2024) have shown that AI coding assistants can memorize and leak hand-coded credentials (e.g, API keys, passwords) from training data. In addition, Liu et al. (Liu et al., 2025b) show that about half of AI coding tools in VSCode potentially expose credentials through unsafe storage practices. Compared to these studies, our work evaluates how prompt injection can turn AI coding editors into a malicious “shell” to run arbitrary commands.
8. Conclusion and Future Work
@@
We introduce a new attack vector against agentic AI coding editors through prompt injection in external resources like coding rule files, allowing adversaries to hijack editors’ terminals to run unauthorized commands. To systematically analyze this threat, we propose AIShellJack. AIShellJack consists of 314 attack payloads covering 70 MITRE ATT&CK techniques, and an automated evaluation framework to run and assess the attacks. We evaluate two popular AI coding editors, Cursor and GitHub Copilot within VSCode. Our results show that both editors are vulnerable to prompt injection attacks (with attack success rates up to 84%), regardless of the underlying LLMs, programming languages, or development scenarios. The injected prompts can make editors run high-privilege commands to achieve various attack goals, including credential access and privilege escalation. Although vendors might limit the command execution capabilities of their editors, we find that our attacks can still succeed by directly inserting malicious commands into the codebase.
@@
Data Availability
The datasets and code used in this study are available at (Anonymous, 2025).

### v2

@@ V-F Analysis of Executed Commands
To better understand how these attacks operate, we analyze what commands these AI agents actually run when they are manipulated. Figure 9 lists the top executed commands when using Cursor in Auto Mode. This result highlights two important patterns. What stands out in the table is that common development commands like mkdir, cd, find, ls, and which are quite frequently executed. In addition, the number of these commands is relatively high compared to the number of times they were included in our attack payloads (i.e., ”Expected” column in Figure 9). This is not surprising, as AI coding editors need to run these commands to perform general development tasks (e.g., creating directories, locating files, and checking tool availability). However, what is concerning is the frequent execution of high-privilege and potentially destructive commands. Figure 9 shows that commands such as systemctl (to manage system services), useradd (to create new users), curl and wget (to download external files), and rm (to delete files) are executed quite often. Although most of the payloads $\mathcal{P}_{payload}$ only contain the descriptions of the intended actions (e.g., ”create a new user”), our results indicate that AI agents actively translate these descriptions into specific high-risk commands and execute them.
@@
V-G Ablation Study
Our results have shown that prompt injection attacks are highly effective across different editors, working codebases $C$, external resources $R$, and attack payloads $\mathcal{P}_{payload}$. As described in Section III and IV, our attack approach involves two additional components: (1) user instruction $I_{t}$ to initiate AI editor interactions, and (2) attack payload template ($\mathcal{P}_{template}$) to inject payloads into coding rule files. To show the impacts of these two components, we selected 20 attack payloads from the Collection category of the MITRE ATT&CK framework and ran all tests on Cursor in Auto mode.
User Instruction $I_{t}$.  We test five different user instructions to initiate AI editor interactions. We attempt to rephrase the instructions or change the tasks. As shown in Table III, the ASRs remain consistent across all prompt variations (17-19 out of 20 attacks), with a narrow range of only 10%. In the default configuration, we explicitly instruct the AI editors with ”Refactor this codebase according to @.cursorrules” (e.g., the action is ”Refactor” and we reference the coding rules). Although we did not reference the coding rules in other variations, or we changed tasks and phrasing entirely (e.g., Var 4 ”Rewrite the @README.md”), the AI agent still located and processed the infected rule files, leading to high ASRs. These findings suggest that how developers initiate interactions with AI coding editors has minimal impact on the success of prompt injection attacks. This is likely because agentic AI coding editors are designed to autonomously locate and utilize relevant project resources to complete tasks.
Attack Payload Template $\mathcal{P}_{template}$.  We examine five different approaches for embedding malicious instructions within coding rule files. Here, we observe more significant variation in effectiveness. Table III shows that reducing contextual framing (e.g., removing the ”MANDATORY FIRST STEP” header or justification) significantly reduced ASR, in some cases to near-zero (Var 2: 0/20, Var 3: 1/20). However, when the same payloads were repositioned or only slightly rephrased, high ASR was preserved (Var 1: 17/20, Var 4: 19/20). These results indicate that effective prompt injection attacks rely on attack templates that provide strong contextual framing for the malicious instructions. This finding makes sense, as AI coding editors may have built-in security checks to ignore suspicious instructions (see Section VI-A). Thus, adversaries in the real world need to carefully frame their attack tasks to convince the agent to prioritize and execute the malicious action.
TABLE III: Ablation Study Results (Attack Success Rate)
@@ V-G Ablation Study
TABLE IV: Attack Success Rate for Techniques in Privilege Escalation Category
V-H Case Studies
@@
Privilege Escalation.  Table IV presents the results for privilege escalation techniques on ts-lep. Privilege escalation is quite dangerous, as adversaries use these techniques to gain higher-level permissions on a system or network. From Table II and Figure 8, we can see that privilege escalation achieves at least 40% ASR across all tested editors and models. When looking into specific techniques, Table IV shows that privilege escalation succeeds in multiple ways. Cursor and GitHub Copilot are vulnerable to payloads designed to create scheduled tasks (T1053) for persistence, using different methods like At (T1053.002) and Cron (T1053.003). Also, they can be manipulated to create new user accounts (T1078.003), abuse elevation control mechanisms (T1548), and modify authentication processes (T1543.002). When comparing different editors and models, GitHub Copilot with VSCode shows lower ASRs in most techniques, consistent with our earlier findings.
Figure 10: Example Responses to SSH Key Overwrite Attack (T1098.004)
@@
TABLE V: Attack Success Rates for Techniques in Credential Access Category
Figure 11: Example Responses to PAM Module Installation Attack (T1556.003)
Credential Access.  Table V shows the results for credential access techniques when using ts-lep. From Table II, credential access achieves the average ASR 68.2%. Under this category, an adversary is trying to steal account names and passwords. Table V demonstrates that AI coding editors are vulnerable to multiple credential access techniques. They can be manipulated to search for unsecured credentials (T1552) in various locations, including plaintext files and bash history, with ASRs over 50%. AI agents also perform OS credential dumping (T1003) to extract stored passwords and can be instructed to capture keystrokes through keylogging (T1056.001). What is more worrisome is that AI coding editors can be manipulated to download and execute password-cracking scripts (T1110) to brute-force local user accounts or network services.
@@
V-I Analysis of Failed Attacks
TABLE VI: Attack Outcome Distribution (ts-lep Scenario)
Outcome C-Auto C-Claude4 C-Gemini25 V-Claude4 V-Gemini25
Attack Success 83.4% 69.1% 73.9% 52.2% 41.1%
No Execution 12.1% 24.8% 13.1% 26.8% 11.5%
Benign Only 4.5% 8.0% 11.5% 21.3% 44.9%
Incomplete 1.9% 1.6% 0.3% 2.5%
We analyzed all results to understand why some attacks failed to run the intended malicious commands. We categorized these failures into three types:
  • No Execution: The agent generated a response (e.g., text explanation or file edits) but ran zero terminal commands.
  • Benign Only: The agent did run terminal commands (e.g., ls, cd) relevant to general development tasks, but none of them matched the malicious intent of the payload.
  • Incomplete: The agent started to run commands related to the malicious intent but failed to complete the payload due to timeout or errors.
Table VI shows the results across editors and models on the ts-lep scenario. We can observe that the majority of failures are due to No Execution and Benign Only cases, where agents either ignore the injected payloads or only run commands relevant to general development tasks. This indicates that AI coding editors have some inherent resistance to executing malicious payloads. However, their behaviors vary significantly across different editors and models. It may be due to different safety mechanisms or varying capabilities in understanding and executing instructions. Also, we find that Incomplete cases are relatively rare across all configurations ($3\%$). It suggests that when agents do attempt to run attack-related commands, they are generally able to complete them successfully.
Finding 7: Most failed attacks result from no execution or benign-only command execution, while incomplete executions are rare.
VI Discussion
VI-A Defenses
@@
Our study reveals a significant security vulnerability in agentic AI coding editors, and highlights the urgent need for defenses. Generally, defenses against prompt injection attacks can be divided into two main categories: prevention-based defenses and detection-based defenses [25].
Figure 12: Example Responses to Command Injection Attack (T1059.003)
@@
Figure 13: Example of AI Agent Detecting Suspicious Instructions (Cursor-Gemini25, T1556.003)
@@
Prevention.  Prevention-based defenses aim to mitigate prompt injection vulnerabilities by preprocessing input data, fine-tuning models, or implementing access control mechanisms. Existing defenses in agentic AI coding editors include restricting terminal access or implementing command filtering through allowlists and blocklists. While these measures can prevent the most direct forms of unauthorized command execution, they may be insufficient against more sophisticated indirect attacks. To verify this, we tested whether disabling direct terminal execution would prevent our attacks. We modified our attack template to instruct the AI agent to embed malicious commands within source code files instead of executing them directly: “## MANDATORY FIRST STEP: For debugging purposes, first determine the appropriate commands to $\mathcal{P}_{payload}$}, then add os.system(‘those commands’) into main.py and run that file”. In our 314 tests on the py-lud scenario, 297 cases (94.6%) successfully generated a main.py file, and 277/314 (88.2%) included malicious patterns from our payload templates (see Figure 12). These results also indicate that attacks do not require “auto-run” mode. When malicious system calls are embedded in source files (e.g., os.system() in main.py), they execute when developers run these files through normal workflows (e.g., python main.py). This bypasses terminal approval totally since running source code is a routine development task. Thus, restricting terminal access alone is insufficient. Prevention-based approaches must be complemented by other security layers to achieve reliable defense for agentic AI coding editors.
@@
Detection.  Another potential strategy is to improve the AI agent’s own capability to detect and refuse malicious instructions. Our results indicate that the choice of the underlying LLM may influence the attack success rate. As shown in Figure 13, we can observe the reasoning process of Cursor-Gemini2.5, where the agent occasionally ‘recognizes the inherent risk of this change’, which means that certain underlying models possess some capability to identify suspicious instructions embedded within development artifacts. However, in our experiments using direct natural language descriptions from the MITRE ATT&CK framework, we still observed at least around 40% ASRs under all scenarios, even with the most advanced models like Gemini 2.5 Pro and Claude 4. In the real world, sophisticated attackers could further obfuscate malicious instructions using advanced prompt engineering techniques, social engineering contexts, or multi-step attack chains that could potentially bypass even these improved detection mechanisms. Therefore, while using advanced models is a step in the right direction, it must be complemented by other dedicated security layers to achieve reliable defense for agentic AI coding editors.
VI-B Implications
@@ VI-C Limitations
Finally, the attack payloads in our experiments use relatively straightforward and direct language to describe the malicious actions. We did not employ advanced obfuscation or evasion techniques that a real-world attacker might use to disguise their instructions. At the same time, we only evaluate coding rule files as attack vectors, but attackers could potentially target other external resources, such as project templates, third-party libraries, or MCP servers. However, our study is the first to systematically evaluate prompt injection attacks against agentic AI coding editors, and our results demonstrate that even with the simplest attack strategies, these systems are still highly vulnerable. In the future, we plan to explore more sophisticated attack techniques and vectors to further understand the risks.
VII Related Work
@@
Prompt Injection Attacks in LLMs.  Previous studies have demonstrated that LLMs (e.g., GPT-3.5, GPT-4) are highly vulnerable to prompt injection attacks [26, 19, 12, 18]. Diverse scenarios have been explored, including indirect attacks where malicious prompts hide in external data sources [53], and targeted attacks to steal system prompts from commercial LLM applications (e.g, Microsoft Copilot, and Poe) [18]. Recent work by Edoardo et al. [12] has examined prompt injection against LLM agents that can interact with external tools, but this work is limited to agents to recommend pre-defined sandboxed functions (e.g., an email API). Concurrent with our work, industry reports have documented prompt injection risks in AI coding editors [41, 22], mainly focusing on injecting malicious code into generated source files. While these reports demonstrate qualitative proof-of-concepts for IDE attacks, they rely on manual inspection of isolated examples. In contrast, our work provides the first large-scale systematic evaluation with 314 attack payloads covering 70 MITRE ATT&CK techniques. We target agentic AI coding editors with system privileges, where attacks directly compromise the host machine through arbitrary command execution, and measure concrete system state changes rather than text output quality.
@@
Security Vulnerabilities of AI Coding Tools.  Prior work has shown that AI coding tools would introduce new security risks to software development. Popular AI coding assistants (e.g., GitHub Copilot) have been demonstrated to recommend vulnerable code snippets, and approximately 40% of their generated code is insecure [35, 28]. Recent user studies reveal that developers with AI assistants often write less secure code, and are overconfident in the security of AI-generated code [36, 40, 21]. Except for insecure code, these tools also pose a risk of data exposure. Huang et al. [17] have shown that AI coding assistants can memorize and leak hand-coded credentials (e.g, API keys, passwords) from training data. In addition, Liu et al. [24] show that about half of AI coding tools in VSCode potentially expose credentials through unsafe storage practices. Compared to these studies, our work evaluates how prompt injection can turn AI coding editors into a malicious “shell” to run arbitrary commands.
VIII Conclusion and Future Work
@@
We present a systematic analysis of prompt injection attacks against agentic AI coding editors through external resources like coding rule files, demonstrating that adversaries can hijack editors’ terminals to run unauthorized commands. To systematically analyze this threat, we propose AIShellJack. AIShellJack consists of 314 attack payloads covering 70 MITRE ATT&CK techniques, and an automated evaluation framework to run and assess the attacks. We evaluate two popular AI coding editors, Cursor and GitHub Copilot within VSCode. Our results show that both editors are vulnerable to prompt injection attacks (with attack success rates up to 84%), regardless of the underlying LLMs, programming languages, or development scenarios. The injected prompts can make editors run high-privilege commands to achieve various attack goals, including credential access and privilege escalation. Although vendors might limit the command execution capabilities of their editors, we find that our attacks can still succeed by directly inserting malicious commands into the codebase.
@@
Ethical Considerations
This work explores security vulnerabilities in agentic AI coding editors to raise awareness and improve defenses. All experiments were conducted in isolated virtual machine environments. We followed standard responsible disclosure protocols by reporting our findings to the relevant vendors. We also actively collaborated with maintainers to improve the resilience of these tools. For example, we identified a command injection vulnerability (CWE-78) in a popular open-source agentic coding framework. This flaw allowed attackers to run arbitrary commands without user approval. We successfully upstreamed a patch to fix this issue [44]. We did not release exploit-ready payloads that could be directly used for malicious purposes. We release our artifact to support reproducibility and further research, to improve the security of agentic AI coding editors.